Please help me with this question!!!! The number of seats in the first 12 rows of a high school auditorium form an arithmetic sequence. The first row has 9 seats. The second row has 11 seats. A) write a recursive formula to represent the sequence. B) write an explicit formula to represent the sequence. C) how many seats are in the 12th row?
Answer by richwmiller(17219)

a) an=a(n-1)+2
b) an=9+(n-1)2
an=9+2n-2
an=2n+7
c) a12=24+7=31
first five 9,11,13,15,17,
rows (6-12) 19,21,23,25,27,29,31